OpenStudy (anonymous):

PLEASE HELP! Algebra 1 standard, 9th grade (PLEASE SHOW WORK, THANK YOU) 4(2-3b)=68

OpenStudy (anonymous):

divide each side by 4

OpenStudy (anonymous):

8-12b=68 -12b=60 b=-5

OpenStudy (anonymous):

thats correct

OpenStudy (anonymous):

THANKS SOOOO MUCH!!!!!! I REALLY APPRECIATE IT!!